OpenStudy (anonymous):

In a right angle triangle, if the two sides are 6 cm, then why is the hypotenuse 6√2?

OpenStudy (inkyvoyd):

 \(a^2+b^2=c^2\) a=b=6 \(\large 6^2+6^2=c^2\) \(\huge 36+36=c^2\) \(\huge c^2=72\) \(\huge c=36*2\) \(\Huge c=6\sqrt{2}\)

OpenStudy (inkyvoyd):

I used the Pythagorean theorem for that.

OpenStudy (anonymous):